# On the Isbell Problem

### v1

@@ II $\tau_{0}(f\upharpoonright 1)$ $\tau_{0}(f\upharpoonright 2)$ $\tau_{0}(f\upharpoonright 3)$ldots
    $D(n,p,\dot{r})=\{a\in\mathbb{Q}\cap\mathcal{M}:(\exists(r,\dot{s})\in D_{n}\cap\mathcal{M})((r,\dot{s})\leq(p,\dot{r})\land r=a)\}\cup\{a\in\mathbb{Q}\cap\mathcal{M}:a\perp p\}$
@@
    For $\mathbb{P}$-names $\dot{r},\dot{s}\in\mathcal{M}$ define $D_{0}(\dot{s},\dot{r})$ as:
    $D_{0}(\dot{s},\dot{r})=\{p\in\mathbb{Q}\cap\mathcal{M}:(\pi(p)\Vdash\dot{r},\dot{s}\in\dot{\mathbb{R}})\land[(\pi(p)\Vdash\dot{r}<\dot{s})\vee(\pi(p)\Vdash\dot{s}<\dot{r})\vee(\pi(p)\Vdash\dot{r}=\dot{s})]\}$
@@
    $D_{1}(p,\dot{r})=\{p_{1}\in\mathbb{Q}\cap\mathcal{M}:(\exists(p_{1},\dot{r}_{1})\in\mathbb{Q}*_{\mathbb{P}}\dot{\mathbb{R}}\cap\mathcal{M})(p_{1}\leq p)\land(p_{1}\Vdash\dot{r}=\dot{r}_{1})\}\cup\{p^{\prime}\in\mathbb{Q}\cap\mathcal{M}:p\perp p^{\prime}\}$
@@
Let $\mathcal{E}$ be the collection of all sets of the types (1)-(4). Let $q_{0}\leq q$ be a $(\mathbb{Q},\mathcal{M},\mathcal{E})$-generic condition. We claim that $q_{0}$ forces each $D_{n}^{\divideontimes}$ to be an open dense subset of ˙$\dot{\mathbb{R}}\cap\mathcal{M}[\dot{G}]$. Let $G$ be a generic filter such that $q_{0}\in G$ and let us work in $V[G]$. First note that ˙$\dot{\mathbb{R}}_{G}=(\mathbb{Q}*_{\mathbb{P}}\dot{\mathbb{R}})^{\divideontimes}$.
@@
First let us see that $(D_{n})^{\divideontimes}$ is dense for each $n\in\omega$. Pick $r\in\dot{\mathbb{R}}_{G}\cap\mathcal{M}[G]$. Then there is $\dot{r}\in\mathcal{M}$ such that $r=\dot{r}_{G}$. Since $D(\dot{r})\cap G\neq\emptyset$, there is $p_{0}\in\mathbb{Q}\cap\mathcal{M}\cap G$ such that $p_{0}\Vdash\dot{r}\in\dot{\mathbb{R}}$. Now, $D_{1}(p_{0},\dot{r})\cap G\neq\emptyset$, and $p_{0}\in G$, so there is $(p_{1},\dot{r}_{1})\in\mathbb{Q}*_{\mathbb{P}}\dot{\mathbb{R}}\cap\mathcal{M}$ such that $p_{1}\Vdash\dot{r}=\dot{r}_{1}$ and $p_{1}\leq p_{0}$. Now, $D(n,(p_{1},\dot{r}_{1}))\cap G\neq\emptyset$, so pick a condition $p_{2}\in D(n,(p_{1},\dot{r}_{1}))\cap G$. Then we should have $p_{2}\leq p_{1}$ and there is $\dot{r}_{2}\in\mathcal{M}$ such that $(p_{2},\dot{r}_{2})\in D_{n}$ and $(p_{2},\dot{r}_{2})\leq(p_{1},\dot{r}_{1})$, which implies $(\dot{r}_{2})_{G}\leq(\dot{r}_{1})=\dot{r}_{G}=r$. By definition of $D_{n}^{\divideontimes}$, we also have $(\dot{r}_{2})_{G}\in D_{n}^{\divideontimes}$. Thus $(D_{n})^{\divideontimes}$ is dense.
@@
Let us now see that $(D_{n})^{\divideontimes}$ is open. Fix $r\in(D_{n})^{\divideontimes}$ and let $s\in\dot{\mathbb{R}}_{G}\cap\mathcal{M}[G]$ be such that $s\leq r$. We have to prove that $s\in D_{n}^{\divideontimes}$. By definition of $D_{n}^{\divideontimes}$, there is $(p_{0},\dot{r})\in D_{n}$ such that $p_{0}\in G$ and $r=\dot{r}_{G}$. Let $\dot{s}\in\mathcal{M}$ be such that $s=\dot{s}_{G}$. Then $D(\dot{s})\cap G\neq\emptyset$. Fix $p_{1}\in D(\dot{s})\cap G$ such that $p_{1}\leq p_{0}$. Then we have $p_{1}\Vdash\dot{s}\in\dot{\mathbb{R}}$. Now, we have $D_{1}(p_{1},\dot{s})\cap G\neq\emptyset$, so we can pick $p_{2}\in D_{1}(p_{1},\dot{s})\cap G$. Note that we have $p_{2}\leq p_{1}$, and there is $\dot{s}_{2}$ such that $(p_{2},\dot{s}_{2})\in\mathbb{Q}*_{\mathbb{P}}\dot{\mathbb{R}}\cap\mathcal{M}$ and $p_{2}\Vdash\dot{s}=\dot{s}_{2}$. Now, since $s\leq r$, there is $p_{3}\in G\cap D(\dot{r},\dot{s}_{2})$ such that $p_{3}\leq p_{2}$ and $\pi(p_{3})\Vdash\dot{s}_{2}\leq\dot{r}$. Note that $(p_{3},\dot{s}_{2})\in\mathbb{Q}*_{\mathbb{P}}\dot{\mathbb{R}}$. Moreover, we also have $(p_{3},\dot{s}_{2})\leq(p_{0},\dot{r})$. Since $(p_{0},\dot{r})\in D_{n}$ and $D_{n}$ is open, we have $(p_{3},\dot{s}_{2})\in D_{n}$, which implies $(\dot{s}_{2})_{G}\in D_{n}^{\divideontimes}$; since $p_{3}\Vdash\dot{s}=\dot{s}_{2}$, we have $s\in D_{n}^{\divideontimes}$. Thus, $D_{n}^{\divideontimes}$ is open. ∎
@@
$(\mathbb{P},\mathbb{Q})$ is $\varsigma$-proper if and only if $\mathbb{Q}$ is strongly proper and $\mathbb{P}$ forces $[\mathbb{Q}:\mathbb{P}]$ is Cohen preserving.
@@
The only if follows from the previous Proposition 6.7 and Lemma 6.8. Let us assume $\mathbb{Q}$ is strongly proper and $\mathbb{P}$ forces $[\mathbb{Q}:\mathbb{P}]$ is Cohen preserving.
Let $\mathcal{M}\prec H(\theta)$ be countable and such that $(\mathbb{P},\mathbb{Q})\in\mathcal{M}$ and pick $q\in\mathbb{Q}\cap\mathcal{M}$. Fix $\dot{\mathbb{R}}\in\mathcal{M}$ a $\mathbb{P}$-name for a partial order. Let $\vec{D}$ be a countable family of open dense subsets of $\mathbb{Q}*_{\mathbb{P}}\dot{\mathbb{R}}\cap\mathcal{M}$. By the previous Lemma 6.9, we can find a $(\mathbb{Q},\mathcal{M})$-generic condition $q^{\prime}\leq q$ forcing each set $D_{n}^{\divideontimes}$ to be an open dense subset of ˙$\dot{\mathbb{R}}\cap\mathcal{M}[\Gamma]$, fix $q_{0}$ one of such conditions. Let $\Gamma_{0}$ be a $\mathbb{P}$-generic filter such that $\pi(q_{0})\in\Gamma_{0}$. Then, in $V[\Gamma_{0}]$, $q_{0}\in[\mathbb{Q}:\mathbb{P}]$ and forces each $D_{n}^{\divideontimes}$ to be an open dense subset of ˙$\dot{\mathbb{R}}\cap\mathcal{M}[\Gamma_{0}]$. Since $[\mathbb{Q}:\mathbb{P}]$ Cohen preserving, by Lemma 6.2, in $V[\Gamma_{0}]$ we can extend $q_{0}$ to a condition $q_{1}$ and find $\{E_{n}:n\in\omega\}$ open dense subsets of ˙$\dot{\mathbb{R}}\cap\mathcal{M}[\Gamma_{0}]$ such that for each $n\in\omega$, $q_{1}\Vdash E_{n}\subseteq D_{n}^{\divideontimes}$. Going back to $V$ we get a $\mathbb{P}$-name $\dot{q}_{1}$ for $q_{1}$ and the corresponding family of $\mathbb{P}$-names $\{\dot{E}_{n}:n\in\omega\}$ for the open dense subsets of ˙$\dot{\mathbb{R}}\cap\mathcal{M}[\Gamma_{0}]$. Then we have $(\pi(q_{0}),\dot{q}_{1})$ forces, for each $n\in\omega$:
@@
    $D_{n}^{\divideontimes}$ is an open dense subset of ˙$\dot{\mathbb{R}}\cap\mathcal{M}[\Gamma_{0}]$.
@@
    If $\mathbb{P}$ is a strongly proper forcing, and $\mathbb{P}$ forces $\dot{\mathbb{Q}}$ to be a strongly proper forcing and Cohen preserving, then $(\mathbb{P},\mathbb{P}*\dot{\mathbb{Q}})$ is $\varsigma$-proper.
@@
    If $\mathbb{P}$ is strongly proper then (,×$(\mathbb{P},\mathbb{P}\times\mathbb{Q}^{\kappa})$ is $\varsigma$-proper, where $\mathbb{Q}^{\kappa}$ may be the countable support product of $\kappa$-many copies of the Sacks forcing, Silver forcing, Miller lite forcing, etc..
Now we come to the central definition of this section.
@@
    $\mathbf{P}_{0}=\langle\mathbb{P}_{\beta},\dot{\mathbb{Q}}_{\beta}^{0}:\beta<\alpha\rangle$ is a countable support iteration.
@@
    For each $\beta<\alpha$, $\mathbb{P}_{\beta}\Vdash\dot{\mathbb{Q}}_{\beta}^{0}\lessdot\dot{\mathbb{Q}}_{\beta}^{1}$, and $\dot{\pi}_{\beta}:\dot{\mathbb{Q}}_{\beta}^{1}\to\dot{\mathbb{Q}}_{\beta}^{0}$ is a $\mathbb{P}_{\beta}$-name for a projection.
@@
For $(p_{0},q_{0}),(p_{1},q_{1})\in\mathbb{P}$, define $(p_{1},q_{1})\leq(p_{0},q_{0})$ if and only if $p_{1}\leq_{\mathbf{P}_{0}}p_{0}$ and for each $\beta\in dom(q_{0})$, $p_{1}\upharpoonright\beta\Vdash q_{1}(\beta)\leq q_{0}(\beta)$.
@@
In the language of partial orders, $(2)$ above is replace by $(2^{\prime})$ below, and we omit the $\mathbb{P}_{\beta}$-name $\dot{\pi}_{\beta}$ (everything else remains the same),
@@
Let $\mathbf{P}$ be a $\varsigma$-iteration and $\Gamma$ a $\mathbf{P}$-generic filter. Let us name all the natural filters arising along a $\varsigma$-iteration. For $\beta<\alpha$ (those with a * will be the most used in the proofs),
  1. 0)*

### v2

@@ II $\tau_{0}(f\upharpoonright 1)$ $\tau_{0}(f\upharpoonright 2)$ $\tau_{0}(f\upharpoonright 3)$ldots
    $\overline{D}(n,p,\dot{r})=\{a\in\mathbb{Q}\cap\mathcal{M}:(\exists(t,\dot{s})\in D_{n})((t,\dot{s})\leq(p,\dot{r})\land t=a)\}$
    Then let $D(n,p,\dot{r})=\overline{D}(n,p,\dot{r})\cup(p^{\perp}\cap\mathcal{M})$
@@
    For $\mathbb{P}$-names $\dot{r},\dot{s}\in\mathcal{M}$. For a condition $p\in\mathbb{Q}$, let us say that $p$ decides the relation $\dot{s}<\dot{r}$ it either, $\pi(p)\Vdash\dot{r}<\dot{s}$, or $\pi\Vdash\dot{r}>\dot{r}$ or $\pi(p)\Vdash\dot{r}=\dot{s}$. define $D_{0}(\dot{s},\dot{r})$ and $D_{1}(\dot{s},\dot{r})$ as:
    $D_{0}(\dot{s},\dot{r})=\{p\in\mathbb{Q}\cap\mathcal{M}:(\pi(p)\Vdash\dot{r},\dot{s}\in\dot{\mathbb{R}})\land(p\text{ decides }\dot{r}<\dot{s}))\}$
@@
    $\overline{D}_{1}(p,\dot{r})=\{p_{1}\in\mathbb{Q}\cap\mathcal{M}:(\exists(p_{1},\dot{r}_{1})\in\mathbb{Q}*_{\mathbb{P}}\dot{\mathbb{R}}\cap\mathcal{M})(p_{1}\leq p)\land(p_{1}\Vdash\dot{r}=\dot{r}_{1})\}$
    and let $D_{1}(p,\dot{r})=\overline{D}_{1}(p,\dot{r})\cup(p^{\perp}\cap\mathcal{B})$.
@@
Let $\mathcal{E}$ be the collection of all sets of the types (1)-(4). Let $q_{0}\leq q$ be a $(\mathbb{Q},\mathcal{M},\mathcal{E})$-generic condition. We claim that $q_{0}$ forces each $D_{n}^{\divideontimes}$ to be an open dense subset of ˙$\dot{\mathbb{R}}\cap\mathcal{M}[\dot{G}]$. Let $G$ be a generic filter such that $q_{0}\in G$ and let us work in $V[G]$. Note that ˙$\dot{\mathbb{R}}_{G}=(\mathbb{Q}*_{\mathbb{P}}\dot{\mathbb{R}})^{\divideontimes}_{G}$.
@@
First let us see that $(D_{n})^{\divideontimes}_{G}$ is dense for each $n\in\omega$. Pick $r\in\dot{\mathbb{R}}_{G}\cap\mathcal{M}[G]$. Then there is $\dot{r}\in\mathcal{M}$ such that $r=\dot{r}_{G}$. Since $D(\dot{r})\cap G\neq\emptyset$, there is $p_{0}\in\mathbb{Q}\cap\mathcal{M}\cap G$ such that $p_{0}\Vdash\dot{r}\in\dot{\mathbb{R}}$. Now, $D_{1}(p_{0},\dot{r})\cap G\neq\emptyset$, and $p_{0}\in G$, so there is $(p_{1},\dot{r}_{1})\in\mathbb{Q}*_{\mathbb{P}}\dot{\mathbb{R}}\cap\mathcal{M}$ such that $p_{1}\Vdash\dot{r}=\dot{r}_{1}$ and $p_{1}\leq p_{0}$. Now, $D(n,(p_{1},\dot{r}_{1}))\cap G\neq\emptyset$, so pick a condition $p_{2}\in D(n,(p_{1},\dot{r}_{1}))\cap G$. Then we should have $p_{2}\leq p_{1}$ and there is $\dot{r}_{2}\in\mathcal{M}$ such that $(p_{2},\dot{r}_{2})\in D_{n}\cap\mathcal{M}$ and $(p_{2},\dot{r}_{2})\leq(p_{1},\dot{r}_{1})$, which implies $(\dot{r}_{2})_{G}\leq(\dot{r}_{1})=\dot{r}_{G}=r$. By definition of $D_{n}^{\divideontimes}$, we also have $(\dot{r}_{2})_{G}\in(D_{n}^{\divideontimes})_{G}$. Thus $(D_{n})^{\divideontimes}_{G}$ is dense.
@@
Let us now see that $(D_{n})^{\divideontimes}$ is open. Fix $r\in(D_{n})^{\divideontimes}_{G}$ and let $s\in\dot{\mathbb{R}}_{G}\cap\mathcal{M}[G]$ be such that $s\leq r$. We have to prove that $s\in(D_{n})^{\divideontimes}_{G}$. By definition of $(D_{n})^{\divideontimes}_{G}$, there is $(p_{0},\dot{r})\in D_{n}$ such that $p_{0}\in G$ and $r=\dot{r}_{G}$. Let $\dot{s}\in\mathcal{M}$ be such that $s=\dot{s}_{G}$. Then $D(\dot{s})\cap G\cap\mathcal{M}\neq\emptyset$. Fix $p_{1}\in D(\dot{s})\cap G\cap\mathcal{M}$ such that $p_{1}\leq p_{0}$. Then we have $p_{1}\Vdash\dot{s}\in\dot{\mathbb{R}}$. Now, we have $D_{1}(p_{1},\dot{s})\cap G\cap\mathcal{M}\neq\emptyset$, so we can pick $p_{2}\in D_{1}(p_{1},\dot{s})\cap G\cap\mathcal{M}$. Note that we have $p_{2}\leq p_{1}$, and there is $\dot{s}_{2}$ such that $(p_{2},\dot{s}_{2})\in\mathbb{Q}*_{\mathbb{P}}\dot{\mathbb{R}}\cap\mathcal{M}$ and $p_{2}\Vdash\dot{s}=\dot{s}_{2}$. Now, since $s\leq r$, there is $p_{3}\in G\cap D(\dot{r},\dot{s}_{2})\cap\mathcal{M}$ such that $p_{3}\leq p_{2}$ and $\pi(p_{3})\Vdash\dot{s}_{2}\leq\dot{r}$. Note that $(p_{3},\dot{s}_{2})\in\mathbb{Q}*_{\mathbb{P}}\dot{\mathbb{R}}$. Moreover, we also have $(p_{3},\dot{s}_{2})\leq(p_{0},\dot{r})$. Since $(p_{0},\dot{r})\in D_{n}$ and $D_{n}$ is open, we have $(p_{3},\dot{s}_{2})\in D_{n}$, which implies $(\dot{s}_{2})_{G}\in D_{n}^{\divideontimes}$; since $p_{3}\Vdash\dot{s}=\dot{s}_{2}$, we have $s\in D_{n}^{\divideontimes}$. Thus, $D_{n}^{\divideontimes}$ is open. ∎
@@
If $\mathbb{Q}$ is $\sigma$-proper and $\mathbb{P}$ forces $[\mathbb{Q}:\mathbb{P}]$ is Cohen preserving, then $(\mathbb{P},\mathbb{Q})$ is Cohen proper.
@@
Let $\mathcal{M}\prec H(\theta)$ be countable and such that $(\mathbb{P},\mathbb{Q})\in\mathcal{M}$ and pick $q\in\mathbb{Q}\cap\mathcal{M}$. Fix $\dot{\mathbb{R}}\in\mathcal{M}$ a $\mathbb{P}$-name for a partial order. Let $\vec{D}$ be a countable family of open dense subsets of $\mathbb{Q}*_{\mathbb{P}}\dot{\mathbb{R}}\cap\mathcal{M}$. By the previous Lemma 6.6, we can find a $(\mathbb{Q},\mathcal{M})$-generic condition $q^{\prime}\leq q$ forcing each set $D_{n}^{\divideontimes}$ to be an open dense subset of ˙$\dot{\mathbb{R}}\cap\mathcal{M}[\dot{\Gamma}]$, fix $q_{0}$ one of such conditions. Let $\Gamma_{0}$ be a $\mathbb{P}$-generic filter such that $\pi(q_{0})\in\Gamma_{0}$. Then, in $V[\Gamma_{0}]$, $q_{0}\in[\mathbb{Q}:\mathbb{P}]$ and forces each $D_{n}^{\divideontimes}$ to be an open dense subset of ˙$\dot{\mathbb{R}}\cap\mathcal{M}[\Gamma_{0}]$. Since $[\mathbb{Q}:\mathbb{P}]$ Cohen preserving, by Lemma 6.2, in $V[\Gamma_{0}]$ we can extend $q_{0}$ to a condition $q_{1}$ and find $\{E_{n}:n\in\omega\}$ open dense subsets of ˙$\dot{\mathbb{R}}\cap\mathcal{M}[\Gamma_{0}]$ such that for each $n\in\omega$, $q_{1}\Vdash E_{n}\subseteq D_{n}^{\divideontimes}$. Going back to $V$ we get a $\mathbb{P}$-name $\dot{q}_{1}$ for $q_{1}$ and the corresponding family of $\mathbb{P}$-names $\{\dot{E}_{n}:n\in\omega\}$ for the open dense subsets of ˙$\dot{\mathbb{R}}\cap\mathcal{M}[\Gamma_{0}]$. Then we have $(\pi(q_{0}),\dot{q}_{1})$ forces, for each $n\in\omega$:
@@
    $D_{n}^{\divideontimes}$ is an open dense subset of ˙$\dot{\mathbb{R}}\cap\mathcal{M}[\dot{\Gamma}_{0}]$.
@@
Theorem 6.8.
Let $\mathbb{P}$ and $\mathbb{Q}$ be forcings such that $\mathbb{P}\lessdot\mathbb{Q}$ and $\pi:\mathbb{Q}\to\mathbb{P}$ a regular projection. The following are equivalent:
  1. (1)
    $(\mathbb{P},\mathbb{Q})$ is Cohen proper.
  2. (2)
    $\mathbb{Q}$ is $\sigma$-proper and [:] $\mathbb{P}\Vdash[\mathbb{Q}:\mathbb{P}]\text{ is Cohen preserving}$.
  3. (3)
    For any $\mathcal{M}\prec H(\theta)$ countable such that $\mathbb{Q},\mathbb{P}\in\mathcal{M}$, any open dense $D\subseteq(\mathbb{Q}\times\mathbb{C})\cap\mathcal{M}$ and any condition $q\in\mathbb{Q}\cap\mathcal{M}$, there are $q_{0}\leq q$ and a $\mathbb{P}$-name $\dot{E}$ such that,
    1. (a)
      $\pi(q_{0})\Vdash\dot{E}\subseteq\mathbb{C}\text{ is open dense}$.
    2. (b)
      $q_{0}\Vdash\dot{E}\subseteq{D}^{\divideontimes}$.
  4. (4)
    For any $\mathcal{M}\prec H(\theta)$ countable such that $\mathbb{Q},\mathbb{P}\in\mathcal{M}$, any $\mathbb{P}$-name $\dot{\mathbb{R}}\in\mathcal{M}$, any open dense $D\subseteq(\mathbb{Q}*_{\mathbb{P}}\dot{\mathbb{R}})\cap\mathcal{M}$ and any condition $q\in\mathbb{Q}\cap\mathcal{M}$, there are $q_{0}\leq q$ and a $\mathbb{P}$-name $\dot{E}$ such that,
    1. (a)
      $\pi(q_{0})\Vdash\dot{E}\subseteq\dot{\mathbb{R}}\cap\mathcal{M}[{\pi[\dot{\Gamma}}]]\text{ is open dense}$.
    2. (b)
      $q_{0}\Vdash\dot{E}\subseteq{D}^{\divideontimes}$.
Proof.
The implications $(4)\Longrightarrow(3)$ and $(1)\Longrightarrow(4)$ are easy, and the implication $(2)\Longrightarrow(1)$ is Proposition 6.7. So we are left with implication $(3)\Longrightarrow(2)$.
[…]
  1. i)
    $q_{0}$ is $(\mathcal{F},\mathbb{Q},\mathcal{M})$-generic.
  2. ii)
    $\pi(q_{0})\Vdash\dot{E}\subseteq\mathbb{C}\text{ is open dense}$.
  3. iii)
    $q_{0}\Vdash\dot{E}\subseteq\dot{D}$.
Let us first define $D_{0}=\{(q,s)\in\mathbb{Q}\times\mathbb{C}:q\Vdash s\in\dot{D}\}\cap\mathcal{M}$. It is clear that $D_{0}$ is an open dense subset of $(\mathbb{Q}\times\mathbb{C})\cap\mathcal{M}$. Define the set $H_{0}$ as follows:
$H_{0}=\{(q,s_{0}^{\frown}t):(q,t)\in D_{0}\}$
Now, for each positive $n\in\omega$, define a set $H_{n}$ as follows:
$(p,t)\in H_{n}\Longleftrightarrow(p,t)\in D_{0}\land p\in F_{n-1}\land s_{n}\subseteq t$
and let $H=\bigcup_{n\in\omega}H_{n}$. Let us see that $H$ is an open dense subset of $\mathbb{Q}\times\mathbb{C}\cap\mathcal{M}$. Let $(r,t)\in\mathbb{Q}\times\mathbb{C}\cap\mathcal{M}$ be arbitrary. We can assume $s_{n}\subseteq t$ for some $n\in\omega$. Since $D_{0}$ is open dense, there is $(p_{0},t_{0})\in D_{0}$ such that $(p_{0},t_{0})\leq(r,t)$. The set $F_{n}$ is open dense, there is $p_{1}\in F_{n}$ such that $p_{1}\leq p_{0}$, and since $D_{0}$ is open, we get $(p_{1},t_{0})\in D_{0}$, so we have $(p_{1},t_{0})\in D_{0}$, $p_{1}\in F_{n}$ and $s_{n}\subseteq t_{0}$, which implies $(p_{1},t_{0})\in H_{n}\subseteq H$, so $H$ is dense. To see that $H$ is open just note that each $H_{n}$ is open. Our hypothesis (3) implies that there is a condition $q_{0}\leq q$ and a $\mathbb{P}$-name $\dot{E}$ such that (3)(a) and (3)(b) hold for $H$. Then we have:
  1. (1)
    $q_{0}\Vdash\dot{E}\subseteq H^{\divideontimes}$.
  2. (2)
    $\pi(q_{0})\Vdash\dot{E}\subseteq\mathbb{C}\text{ is open dense}$.
Since $q_{0}$ forces $H^{\divideontimes}$ to contain a dense open set, we have that each $F_{n}$ is predense below $q_{0}$, since otherwise there is $n\in\omega$ such that $H^{\divideontimes}$ is not dense below $s_{n}$. On the other hand, $H_{0}^{\divideontimes}$ is open dense below $s_{0}$, and since $H_{0}$ is isomorphic to $D_{0}$, we have that $D_{0}^{\divideontimes}$ is open dense in $2^{<\omega}$. Finally note that $q_{0}\Vdash D_{0}^{\divideontimes}\subseteq\dot{D}_{\Gamma}$: let $G$ be a generic filter such that $q_{0}\in G$ and pick $t\in D_{0}^{\divideontimes}$, so there is $p\in G$ such that $(p,t)\in D_{0}$, and by definition of $D_{0}$ we have $p\Vdash t\in\dot{D}$, so it follows $t\in\dot{D}_{G}$. ∎
@@
    If $\mathcal{U}$ is a suitable filter, then $(\mathbb{S},\mathbb{S}\times\mathbb{P}_{\kappa}(\mathcal{U}))$ is Cohen proper.
@@
    If $\mathbb{P}$ is a $\sigma$-proper forcing, and $\mathbb{P}$ forces $\dot{\mathbb{Q}}$ to be a $\sigma$-proper forcing and Cohen preserving, then $(\mathbb{P},\mathbb{P}*\dot{\mathbb{Q}})$ is Cohen proper.
  3. (3)
    If $\mathbb{P}$ is $\sigma$-proper then (,×$(\mathbb{P},\mathbb{P}\times\mathbb{Q}^{\kappa})$ is Cohen proper, where $\mathbb{Q}^{\kappa}$ may be the countable support product of $\kappa$-many copies of the Sacks forcing, Silver forcing, Miller lite forcing, etc.
7. Iteration of Cohen properness.
Here we define a way to iterate Cohen proper forcings which seems very similar to countable support iterations and allows us to get the continuum arbitrarily large.
@@
    $\mathbf{P}_{0}^{\alpha}=\langle\mathbb{P}_{\beta},\dot{\mathbb{Q}}_{\beta}^{0}:\beta<\alpha\rangle$ is a countable support iteration.
@@
    For each $\beta<\alpha$, $\mathbb{P}_{\beta}\Vdash\dot{\mathbb{Q}}_{\beta}^{0}\lessdot\dot{\mathbb{Q}}_{\beta}^{1}$, and $\dot{\pi}_{\beta}:\dot{\mathbb{Q}}_{\beta}^{1}\to\dot{\mathbb{Q}}_{\beta}^{0}$ is a $\mathbb{P}_{\beta}$-name for the regular projection.
@@
For $(p_{0},q_{0}),(p_{1},q_{1})\in\mathbf{P}$, define $(p_{1},q_{1})\leq(p_{0},q_{0})$ if and only if $p_{1}\leq_{\mathbf{P}_{0}}p_{0}$ and for each $\beta\in dom(q_{0})$, $p_{1}\upharpoonright\beta\Vdash(p_{1}(\beta),q_{1}(\beta))\leq(p_{0}(\beta),q_{0}(\beta))$.
@@
In the language of partial orders, $(2)$ above is replaced by $(2^{\prime})$ below, and we omit the $\mathbb{P}_{\beta}$-name $\dot{\pi}_{\beta}$ (everything else remains the same),
@@
Let $\mathbf{P}=\langle\mathbb{P}_{\beta},(\dot{\mathbb{Q}}_{\beta}^{0},\dot{\mathbb{Q}}_{\beta}^{1},\dot{\pi}_{\beta}):\beta<\alpha\rangle$ be a $\varsigma$-iteration. We say that $\mathbf{P}$ is a Cohen proper $\varsigma$-iteration, if:
  1. (1)
    $(\mathbb{Q}_{0}^{0},\mathbb{Q}_{0}^{1})$ is Cohen proper.
  2. (2)
    For all $\beta<\alpha$, $\mathbb{P}_{\beta}\Vdash(\dot{\mathbb{Q}}_{\beta}^{0},\dot{\mathbb{Q}}_{\beta}^{1})$ is Cohen proper.
Definition 7.3.
Along a $\varsigma$-iteration there are many different generic filters arising, and we name some of them we will need later. For $\beta<\alpha$,
  1. 0)